OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)). 23 47 82 90

OpenStudy (anonymous):

What do you think it is

OpenStudy (anonymous):

47?

OpenStudy (anonymous):

One sec let me check my answer

OpenStudy (anonymous):

ok

geerky42 (geerky42):

Firstly, what is g(3)?

OpenStudy (anonymous):

I don't know. That's why I'm asking for help obviously..

geerky42 (geerky42):

Well, all you have to do is replace x to 3.

OpenStudy (anonymous):

Its 47

OpenStudy (er.mohd.amir):

find g(3) put 3 at place of x in g(x).

OpenStudy (anonymous):

Thank you @bleebleebleed

geerky42 (geerky42):

\[g(x) = x^2+1\\~\\g(\color{red}{3}) = \color{red}{3}^2+1\]

OpenStudy (anonymous):

@emilyfuchs your welcome

geerky42 (geerky42):

Are you sure it's 47? lol...

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

POny!! Save them D: !

geerky42 (geerky42):

Work isn't hard to do really... What is g(3)?

geerky42 (geerky42):

\[g(3)= 3^2 + 1 =~?\]

OpenStudy (anonymous):

10?

geerky42 (geerky42):

Yeah, so \[f(g(3)) = f(10)\] Now find f(10)

OpenStudy (anonymous):

S o is the answer 23?

geerky42 (geerky42):

\[f(x) = 2x+3\\~\\ f(\color{red}{10}) = 2\cdot\color{red}{10}+3=~?\]

geerky42 (geerky42):

Correct.

geerky42 (geerky42):



OpenStudy (anonymous):

@geerky42 LOL Now do the Nae Nae

OpenStudy (anonymous):